solve:
4b -3 (7b + 2) = b - 9

Answer by checkley77(12844)

4b -3 (7b + 2) = b - 9
4b-21b-6=b-9
-17b-b=-9+6
-18b=-3
b=-3/-18
b=1/6
proof:
4*1/6-3(7*1/6+2)=1/6-9
4/6-3(7/6+2)=-53/6
4/6-3(19/6)=-53/6
4/6-57/6=-53/6
-53/6=-53/6